OpenStudy (anonymous):

When x is divided by 7 the remainder is 2. What is the remainder when x+5 is divided by 7?

OpenStudy (anonymous):

lets try an example if you divide 23 by 7 the remainder is 2 what is the remainder when you divide 23+ 5= 28 by 7 ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Is that because 5+2=7, leaving no remainder?

OpenStudy (anonymous):

@satellite73 So if you add 6 instead, the remainder would be 1?

OpenStudy (phi):

yes. you could say x/7 = n (some integer), plus remainder 2 that means 7n+2= x (undo the division) now add 5 to both sides 7n+2+5= x+5 or 7n+7 = x+5 or 7(n+1)= x+5 and the left side is evenly divisible by 7 no remainder if we added 6 instead of 5 7n+8 = x+6 and dividing the left side by 7: (7n+8)/7 = 7n/7 +8/7 = n+1 remainder 1